OpenStudy (anonymous):

Another step-by-step PLEASE!!! This time it's a word problem: A colony of bacteria is growing exponentially according to the function B(t)= 48e0.8t where t is in hours. How many bacteria are there after 8 hours?

OpenStudy (anonymous):

just substitute 8 for t...and you'd be good to go!

OpenStudy (anonymous):

Jenn.. t is number of hours.. so plug in t=8 in the equation and solve..

OpenStudy (anonymous):

B(t)=4*e^0.8t I get confused with the exponent and the t being on BOTH sides of the equal sign. can you please walk me through it?

OpenStudy (anonymous):

ahh.. the first t is a notation.. Like b(t) = b(9)

OpenStudy (anonymous):

so it's kind of like the F(x) where it means Function?

OpenStudy (anonymous):

B(8) = 48 e ^ (0.8*8) = 48 e^(6.4) Now you are good to go...

OpenStudy (anonymous):

Yes exactly.. instead of f(x) it is b(t)

OpenStudy (anonymous):

if that's the case, then I need to plug in 8...let me work it out, and explain it back to you. please correct my thinking as I go

OpenStudy (anonymous):

Sure..

OpenStudy (anonymous):

e^6.4 = 601.8450379 then I multiply it by 4? to get 2407.380151 yes? comments?

OpenStudy (anonymous):

If you are interested in getting tutoring then contact me at helpingtutors@gmail.com

OpenStudy (anonymous):

Is it not 48?

OpenStudy (anonymous):

just need to know if I got it right at this time please!

OpenStudy (anonymous):

It's 4

OpenStudy (anonymous):

Then.. you are right!! your answer is perfect...

OpenStudy (anonymous):

excellent! Thanks!

OpenStudy (anonymous):

But you cannot say number of bacterias in decimal, so round it off to the nearest integer